The Saunders Investment Bank has the following financing outstanding.

Debt:

120,000 bonds with a coupon rate of 8 percent and a current price quote of 110; the bonds have 20 years to maturity. 290,000 zero coupon bonds with a price quote of 17.5 and 30 years until maturity. Assume semiannual compounding.

Preferred stock:

210,000 shares of 6 percent preferred stock with a current price of $70, and a par value of $100.

Common stock:

3,200,000 shares of common stock; the current price is $56, and the beta of the stock is 1.05.

Market:

The corporate tax rate is 40 percent, the market risk premium is 7 percent, and the risk-free rate is 4 percent.

What is the WACC for the company? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

Solve for WACC in percentage.
Answer #1

SOLUTION:

calculation of market value weight

Financing

working

market value

Weight

8% bond

120000*1100

132000000

35.05%

zero coupon bond

290000*175

50750000

13.47%

preferred stock

210000*70

14700000

3.90%

Equity

3200000*56

179200000

47.58%

total market value

376650000

100.00%

Calculation of cost of capital:

8 percent bond:

1100=40*(PVIFA r%,40) + 1000*(PVIF r%,40)

r=3.53%

since the coupon payments are semi-annual , YTM on bonds is

3.53%*2

7.06%

after tax cost of 8% coupon debt = 7.06*(1-0.4)

4.24%

zero coupon bond:

175=1000*(PVIF r%,60)

r=2.95%

since the coupon payments are semi-annual , YTM on bonds is

2.95%*2

5.90%

after tax cost of zero coupon debt = 5.90*(1-0.4)

3.54%

cost of preferred stock (rp)=Dividend/market price=6/70

8.57%

cost of equity (re) = risk free rate + beta*market risk premium

0.04+1.05*0.07

11.35%

calculation of WACC

Financing

after tax cost of capital

weight

a

b

a*b

8% bond

4.24%

35.05%

1.48%

zero coupon bond

3.54%

13.47%

0.48%

preferred stock

8.57%

3.90%

0.33%

Equity

11.35%

47.58%

5.40%

WACC

7.70%